# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 8.1 The Attention-Recurrence Expressivity Gap
Still, the work by Jelassi et al. [124] showed that these mechanisms do not seem to be sufficient to close the recurrence expressivity gap. To counteract this, the Griffin model includes a local attention [17], also known as sliding window attention, into SSM-based models (i.e., LRU [213]). In practice, local attention allows each position to attend only to a fixed number of tokens in the past, reducing the number of FLOPs with respect to standard attention and bounding the size of the KV-cache to the size of window, making it no longer quadratic in the sequence length. Griffin matches the learning speed of Transformers (in terms of necessary training steps/samples to achieve a certain performance), and is capable to extrapolate to evaluation sequences several orders of magnitude longer than the training sequence lengths.
These findings show that recurrent architectures working in the state-space have still room for improvement, in terms or representational capability, when compared to attention-based methods. Future work should focus on surpassing the current approach to build hybrid architectures that endow state-space models with an attention-like mechanism and investigate other solutions. Moreover, it is of extreme interest to understand the role of model pre-training. Indeed, Amos et al. [3] suggested that self-supervised pretraining on the task data (referred to as self pretraining), rather than training from scratch, allows to strongly increase the performances on long sequences (i.e., on LRA) both in vanilla Transformers and with simple diagonal RNNs, showing the redundancy of sophisticated architectures and manually-designed biases [87, 171]
@@
To address the low FLOPs-to-byte ratio due to elementwise operations, the Griffin model [57] is implemented in TPUs and leverages a custom kernel for the computation of the state update equation exploiting linear scans. The implemented kernel minimizes memory transfers by keeping hidden states in the TPUs Vector Memory (VMEM) instead of HBM and performing memory transfers in larger chunks, resulting in a 3×3$\times$3 × speedup with respect to a vanilla implementation. As an interesting remark, the gating mechanism in Griffin is not compatible with the usual convolutional view or associative scans proper of LTI SSMs. Even if the authors could have used associative scans similarly to GateLoop [135], they found that these operations reduce the number of FLOPs, but do not reduce memory overheads, which is their primary bottleneck.
Summing up, when dealing with long-sequence processing, a very relevant role is played by the way computations are optimized, and identifying primary performance bottlenecks on the target-hardware/modern-accelerators represent a fundamental step for future research. Of course, novel models in this research field could highly benefit from ad-hoc hardware and algorithmic solutions.
@@ 8.3 Infinite-length Sequences: Lifelong Learning
This survey describes multiple architectural and algorithmic solutions that allow neural networks to handle sequences of increasing length, by reducing the overall computational complexities and leveraging a stateful recurrent representation. In the previous subsections, we showcased two very relevant research avenues for future studies, i.e., the expressivity improvement of recurrence-based models with respect to full-fledged attention, and the need for hardware-aware solutions capable of effectively dealing with very long sequences.
@@
However, there is still a significant gap that requires specific studies to move beyond current technologies, that shows up when thinking of transitioning from long sequences to infinite-length sequences. Indeed, while researchers aim to mimic human learning abilities from sequential data, the most common existing approaches primarily deal with finite-length datasets BPTT in stateful models, or standard Back-Propagation, whilst the alternative learning algorithms we described in Section 6 are still not widespread in the literature. This stands in stark contrast to human learning from continuous perceptual stimuli, which unfold over time without being pre-segmented into finite-length sequences for stochastic gradient descent optimization, and without the need to explicitly store exemplars into buffers or the of need building data collections [79]. Moreover, this misalignment deviates from the principles of seminal works in recurrent models. Indeed, the goal of works such as the one by Williams and Peng [302] was to devise “an online algorithm, designed to be used to train a network while it runs; no manual state resets or segmentations of the training stream is required”. Even the original formulations of LSTMs, among others [105], were introduced with a learning algorithm that unlike BPTT is “local in space and time”, where “there is no need to store activation values observed during sequence processing in a stack with potentially unlimited size”.
@@
In contrast, as we described in the previous Sections, the most popular research activities in the field of RNNs are not considering the case of learning online from a continuous, possibly infinite, stream of data (i.e., infinite-length sequence), without resets or segmentations [19]. This extremely challenging scenario involves two primary demands: ($i.$) computing meaningful gradients at time $t$ based on processed data up to that point, without fully unrolling the network across the entire (potentially infinite) input sequence (in some cases, only relying on information from times $t$ and $t-1$), and updating model parameters at each time step; ($ii.$) possessing the ability to adapt to the characteristics of the streamed data, retaining acquired skills over time, as emphasized in continual learning contexts [216]. In Section 6 we described the scientific efforts aimed at developing alternatives to BPTT. We showcased online learning algorithms for RNNs, particularly addressing requirement ($i.$). RTRL stands as the most recognized method for online learning with recurrent models, which has been recently refined to mitigate computational and memory constraints [117], alongside various direct or indirect adaptations [184]. We described recent studies that explore forward computation of gradients, enabled by state-space modeling [320], albeit reliant on finite-length sequences processed without updating learnable parameters at each time step, thereby not fully satisfying requirement ($i.$). Considering demand ($ii.$) leads to a scenario akin to Online Continual Learning [216] with substantial overlap with Streaming Learning, as outlined in [93]. It’s important to distinguish this from approaches focused solely on continual learning from streams of finite-length sequences [49, 62].
@@
Summing up, the scenario described in this section remains largely unexplored. Despite advancements in related domains, meeting requirements ($i.$) and ($ii.$) independently poses significant challenges, and their combined fulfillment presents an even more daunting task. Nonetheless, this scenario reflects the reality of learning from continuous streams of visual data, auditory stimuli, perpetual video feeds, language inputs, etc. Whilst common benchmarks for sequential data (see Section 7) do provide valuable opportunities to assess the fulfillment of requirement ($i.$), it is extremely needed to operate within the typical framework of online learning problems, where all data contributes to training and errors accumulate during learning. [^31] The input at time $u_{t}$ is processed and the produced output $y_{t}$ is compared with the available ground truth to update the accuracy estimate; then, $y_{t}$ is used to evaluate a loss function, compute gradients at $t$ and update the model. Certain tasks such as time-series forecasting offer greater flexibility in this context, compared to classification tasks, due to the fact that they provide ground truth at each $t$. Additionally, the properties of the series might change over time, partially allowing to explore requirement ($ii.$). However, a comprehensive evaluation of property ($ii.$) remains challenging with existing datasets.
@@ 9 Conclusions
While Transformers initially overshadowed Recurrent Neural Networks, recent developments in large-context Transformers and (deep) State-Space Models have highlighted the importance of recurrent computations as a viable road to go beyond the limitations of Transformers both in terms of performances on long sequences and scalability. This survey outlined the latest trends and approaches for sequential data processing, offering insights about recent architectural and novel algorithmic solutions. Interestingly, the recent literature in the context of recurrent long-context Transformers and of State-Space Models share the same features, leading towards a computational model based on stateful computations, confirming the advantages brought by the idea of working in the state space. We surveyed the most important works, and we discussed novel trends in learning algorithms for sequential data, offering a broad and detailed perspective on the last generation of neural architectures. This survey also outlined novel research opportunities. In fact, we emphasized how the research direction of lifelong online learning from a stream of data, intended to be an infinite-length sequence, represents a challenge that is still very open and might be faced starting from this renewed interest in recurrent models.

### v2

@@ 8.1 Limitations of Transformers and Recurrent Models
These findings show that recurrent architectures working in the state-space still have room for improvement in terms of representational capability when compared to attention-based methods. Merril et al. [311] showed that, similarly to Transformers, RNNs with a diagonal transition matrix (thus, most of deep SSMs) could only represent functions in TC0. They also showed that the expressive power of SSMs can be extended by making the state transition matrices input-dependent, enabling the theoretical capability to solve NC1-hard problems and achieving better empirical results, as evidenced by models like Liquid-S4 [299]. Grazzi et al. [312] demonstrate that typical linear RNNs like Mamba and DeltaNet are limited by positive-only eigenvalues in their state-transition matrices, which prevents them from solving simple tasks such as parity. By allowing negative eigenvalues and involving non-diagonal state transition matrices, they significantly improve performance on state-tracking, while maintaining stability and efficiency at scale. Concurrently, RWKV-7 [313] demonstrates remarkable expressivity, with the ability to recognize all regular languages using a small constant number of layers. This is grounded in the architecture’s generalized delta rule, which incorporates vector-valued gating and in-context learning rates. These advancements build upon and extend the capabilities introduced in Gated DeltaNet [314], a linear recurrent neural network architecture that combines the strengths of Mamba2’s gating mechanism for adaptive memory decay and DeltaNet’s delta update rule for precise key-value memory updates. Building on [312], Siems et al. [315] propose DeltaProduct, a generalization of DeltaNet where each token update applies a product of multiple Householder transformations. This increases the rank of the state-transition matrix, offering a tunable balance between expressivity and efficiency.
Overcoming limitations with hybrid models.   The expressivity gap emphasized by Jelassi et al. [241] highlights why SSM-based architectures still lag behind attention-only models like Mistral 7B [316] in demanding tasks such as language modeling. To address this, recent models like Griffin [44] incorporate hybrid strategies, blending SSMs with local (sliding window) attention. This form of attention enables each token to access a limited context of past tokens, reducing FLOPs and constraining KV-cache sizes. In practice, local attention allows each position to attend only to a fixed number of tokens in the past, reducing the number of FLOPs with respect to standard attention and bounding the size of the KV-cache to the size of window, making it no longer quadratic in the sequence length. Griffin matches the learning speed of Transformers (in terms of necessary training steps/samples to achieve a certain performance), and is capable to extrapolate to evaluation sequences several orders of magnitude longer than the training sequence lengths. Other architectures have adopted different hybrid schemes, layering attention and SSMs in alternation. For instance, Jamba and Jamba 1.5 [317] propose a Transformer-Mamba mix, combined with a mixture-of-experts module, to achieve a balance between performance and computational cost. Poli et al. [318] conducted a comprehensive analysis and found that a compute-optimal configuration includes roughly one-fourth self-attention layers, with the remainder composed of SSMs. Zamba [319], a 7B hybrid model, demonstrates strong performance relative to other open-weight models by pairing a Mamba backbone with a shared attention layer, thereby capturing the benefits of attention while keeping parameter usage minimal.
Moreover, it is of extreme interest to understand the role of model pre-training. Indeed, Amos et al. [320] suggested that self-supervised pretraining on the task data (referred to as self pretraining), rather than training from scratch, significantly improves the performance on long sequences (i.e., on LRA) both in vanilla Transformers and with simple diagonal RNNs, showing the redundancy of sophisticated architectures and manually-designed biases [150, 135].
@@
To address the low FLOPs-to-byte ratio due to elementwise operations, the Griffin model [44] is implemented in TPUs and leverages a custom kernel for the computation of the state update equation exploiting linear scans. The implemented kernel minimizes memory transfers by keeping hidden states in the TPUs Vector Memory (VMEM) instead of HBM and performing memory transfers in larger chunks, resulting in a $3\times$ speedup with respect to a vanilla implementation. As an interesting remark, the gating mechanism in Griffin is not compatible with the usual convolutional view or associative scans proper of LTI SSMs. Even if the authors could have used associative scans similarly to GateLoop [143], they found that these operations reduce the number of FLOPs, but do not reduce memory overheads, which is their primary bottleneck. Summing up, when dealing with long-sequence processing, a very relevant role is played by the way computations are optimized, and identifying primary performance bottlenecks on the target hardware or modern accelerators represents a fundamental step for future research. Novel models in this research field could greatly benefit from ad-hoc hardware and algorithmic solutions.
@@ 8.3 Infinite-length Sequences: Lifelong Learning
This survey describes multiple architectural and algorithmic solutions that allow neural networks to handle sequences of increasing length, by reducing the overall computational complexities and leveraging a stateful recurrent representation. In the previous subsections, we showcased two very relevant research avenues for future studies, i.e., the expressivity improvement of recurrence-based models with respect to full-fledged attention, and the need for hardware-aware solutions capable of effectively dealing with very long sequences. However, there is still a significant gap that requires specific studies to move beyond current technologies, that shows up when thinking of transitioning from long sequences to infinite-length sequences. Indeed, while researchers aim to mimic human learning abilities from sequential data, the most common existing approaches primarily deal with finite-length datasets BPTT in stateful models, or standard Back-Propagation, whilst the alternative learning algorithms we described in Section 6 are still not widespread in the literature. This stands in stark contrast to human learning from continuous perceptual stimuli, which unfold over time without being pre-segmented into finite-length sequences for stochastic gradient descent optimization, and without the need to explicitly store exemplars into buffers or the need of building data collections [51]. Moreover, this misalignment deviates from the principles of seminal works in recurrent models. Indeed, the goal of works such as the one by Williams and Peng [20] was to devise “an online algorithm, designed to be used to train a network while it runs; no manual state resets or segmentations of the training stream are required”. Even the original formulations of LSTMs, among others [28], were introduced with a learning algorithm that unlike BPTT is “local in space and time”, where “there is no need to store activation values observed during sequence processing in a stack with potentially unlimited size”.
@@
In contrast, as we described in the previous Sections, the most popular research activities in the field of RNNs are not considering the case of learning online from a continuous, possibly infinite, stream of data (i.e., infinite-length sequence), without resets or segmentations [327]. This extremely challenging scenario involves two primary demands: ($i.$) computing meaningful gradients at time $t$ based on processed data up to that point, without fully unrolling the network across the entire (potentially infinite) input sequence (in some cases, only relying on information from times $t$ and t1t-1), and updating model parameters at each time step; (ii.ii.) possessing the ability to adapt to the characteristics of the streamed data, retaining acquired skills over time, as emphasized in continual learning contexts [70]. In Section 6 we described the scientific efforts aimed at developing alternatives to BPTT. We showcased online learning algorithms for RNNs, particularly addressing requirement ($i.$). RTRL stands as the most recognized method for online learning with recurrent models, which has been recently refined to mitigate computational and memory constraints [207], alongside various direct or indirect adaptations [45]. We described recent studies that explore forward computation of gradients, enabled by state-space modeling [212], albeit reliant on finite-length sequences processed without updating learnable parameters at each time step, thereby not fully satisfying requirement ($i.$). Considering demand (ii.ii.) leads to a scenario akin to Online Continual Learning [70] with substantial overlap with Streaming Learning, as outlined in [193]. It’s important to distinguish this from approaches focused solely on continual learning from streams of finite-length sequences [69, 9].
@@
Summing up, the scenario described in this section remains largely unexplored. Despite advancements in related domains, meeting requirements ($i.$) and (ii.ii.) independently poses significant challenges, and their combined fulfillment presents an even more daunting task. Nonetheless, this scenario reflects the reality of learning from continuous streams of visual data, auditory stimuli, perpetual video feeds, language inputs, and more. Whilst common benchmarks for sequential data (see Section 7) do provide valuable opportunities to assess the fulfillment of requirement ($i.$), it is crucial to operate within the typical framework of online learning problems, where all data contributes to training and errors accumulate during learning. [^33] The input at time $u_{t}$ is processed and the produced output $y_{t}$ is compared with the available ground truth to update the accuracy estimate; then, $y_{t}$ is used to evaluate a loss function, compute gradients at $t$ and update the model. Certain tasks such as time-series forecasting offer greater flexibility in this context, compared to classification tasks, due to the fact that they provide ground truth at each $t$. Additionally, the properties of the series might change over time, partially allowing exploration of requirement (ii.ii.). However, a comprehensive evaluation of property (ii.ii.) remains challenging with existing datasets.
@@ 9 Conclusions
While Transformers initially overshadowed Recurrent Neural Networks, recent developments in large-context Transformers and (deep) State-Space Models have highlighted the importance of recurrent computations as a viable road to go beyond the limitations of Transformers both in terms of performance on long sequences and scalability. This survey outlines the latest trends and approaches for sequential data processing, offering insights into recent architectural and novel algorithmic solutions.
Interestingly, the recent literature in the context of recurrent long-context Transformers and State-Space Models share the same features, leading towards a computational model based on stateful computations, confirming the advantages brought by the idea of working in the state space. We surveyed the most important works, and we discussed novel trends in learning algorithms for sequential data, offering a broad and detailed perspective on the last generation of neural architectures. This survey also outlined novel research opportunities. In fact, we emphasized how the research direction of lifelong online learning from a stream of data, intended to be an infinite-length sequence, represents a challenge that is still very open and might be faced starting from this renewed interest in recurrent models.